### v1

# On Prior Distributions for Orthogonal Function Sequences [^1] August 21, 2025

Abstract
We propose a novel class of prior distributions for sequences of orthogonal functions, which are frequently required in various statistical models such as functional principal component analysis (FPCA). Our approach constructs priors sequentially by imposing adaptive orthogonality constraints through a hierarchical formulation of conditionally normal distributions. The orthogonality is controlled via hyperparameters, allowing for flexible trade-offs between exactness and smoothness, which can be learned from the observed data. We illustrate the properties of the proposed prior and show that it leads to nearly orthogonal posterior estimates. The proposed prior is employed in Bayesian FPCA, providing more interpretable principal functions and efficient low-rank representations. Through simulation studies and analysis of human mobility data in Tokyo, we demonstrate the superior performance of our approach in inducing orthogonality and improving functional component estimation.
Key words: Basis function expansion; Functional principal component analysis; Gibbs sampler; Markov Chain Monte Carlo
Introduction
Modeling a sequence of orthogonal functions is required in a variety of statistical problems. Examples include principal component analysis (PCA) for functional data (Margaritella et al., 2021; Suarez and Ghosal, 2017), interactive fixed effects for panel data (Bai, 2009) and Karhunen-Loève expansion for stochastic processes (Yao et al., 2005). Such models entail a sequence of orthogonal functions to achieve interpretable and efficient representations. Although a Bayesian framework needs to specify a prior distribution over the sequence of functions, introducing priors that respect the orthogonality constraint is nontrivial. For instance, existing Bayesian FPCA typically place priors without explicitly enforcing orthogonality (e.g., Suarez and Ghosal, 2017). While such formulations offer computational convenience, they may compromise interpretability and the efficiency of representation, which are central motivations for modeling with orthogonal functions in the first place.
There is a rich literature on incorporating constraints into Bayesian models, particularly in the context of shape-constrained function estimation. For example, monotonicity and convexity constraints have been widely studied in Bayesian nonparametric regression (Shively et al., 2009; Lin and Dunson, 2014; Lenk and Choi, 2017), along with other structural constraints such as unimodality and bounded variation (Wheeler et al., 2017; Yu et al., 2023). While these works focus on constraints imposed on single functions, enforcing mutual orthogonality among multiple functions poses additional challenges due to the rapidly increasing number of pairwise constraints. Several approaches have been proposed to encourage diversity among function components. Plumlee and Joseph (2018) proposed constraint Gaussian process to ensure orthogonality of a single function and pre-specified multiple functions. Constraint relaxation methods (Duan et al., 2020; Matuk et al., 2022) introduce soft penalties to approximate orthogonality, but these typically rely on a fixed tuning parameter, which may require careful calibration and cannot be adaptively learned from the data. As related approaches, repulsive priors (Petralia et al., 2012; Xie and Xu, 2020) can promote separation between functions but do not directly ensure orthogonality in the function space. Other alternatives include orthogonal random Fourier features (Yu et al., 2016), which produce random orthogonal basis functions but do not preserve orthogonality when used to construct multiple functional components. Similarly, models defined on the Stiefel manifold (e.g., Jauch et al., 2025) provide a principled way to impose orthogonal constraints through matrix-valued priors, though these can be computationally demanding and are often impractical when the basis functions themselves are not orthogonal. Post-hoc approaches that modify posterior samples to satisfy constraints (e.g., Sen et al., 2018) offer a practical alternative, but still lack theoretical guarantees.
In this work, we propose the adaptive orthogonal prior (AOP), a novel prior distribution that directly induces approximate orthogonality among functions sequentially, in a computationally efficient and flexible manner. Our paper builds upon the framework of basis function expansions, where each function is represented as a linear combination of pre-specified basis functions. This formulation provides a flexible and interpretable structure that has been widely adopted in Bayesian modeling (e.g. Lang and Brezger, 2004) and offers a unified representation of functions, whose infinite-dimensional limit is known as Gaussian processes (Kimeldorf and Wahba, 1970; Williams and Rasmussen, 2006). Based on the basis function representation, AOP sequentially defines the prior distributions for the coefficient vectors of each function using conditionally Gaussian distributions with approximate orthogonality constraints. By controlling the strength of these constraints through hyperparameters, our prior flexibly accommodates both exact and approximate orthogonality, allowing researchers to balance strictness and model complexity depending on the application. AOP can be seamlessly incorporated into Bayesian functional principal component analysis (FPCA), where orthogonality among principal component functions is crucial for both interpretability and dimensionality reduction. Our prior not only guarantees nearly orthogonal posterior estimates of the principal functions but also enables automatic selection of the effective number of components owing to the sequential construction of AOP, leading to parsimonious low-rank representations. Furthermore, the conditional Gaussian formulation allows for straightforward implementation using standard Gibbs sampling, making posterior inference computationally tractable even in higher-dimensional settings.
This paper is organized as follows. In Section 2, we introduce the AOP and discuss its theoretical properties. Section 3 presents the Bayesian FPCA model incorporating the proposed prior. We demonstrate the numerical performance of the proposed method through simulation studies in Section 4 and an application to human mobility data in Tokyo in Section 5. Finally, concluding remarks are given in Section 6. The R code implementing the proposed method is available at GitHub repository (https://github.com/sshonosuke/AOP).
Prior for Orthogonal Function Sequences
Prior construction
We wish to construct priors for a set of orthogonal functions, $f_{1}(x),\ldots,f_{K}(x)$. To this end, we consider basis function expansions known as a flexible nonparametric approach for function estimation (e.g. Lang and Brezger, 2004). For example, Gaussian process models can be viewed as the infinite-dimensional limit of basis function models (Kimeldorf and Wahba, 1970; Williams and Rasmussen, 2006), where prior on functions is induced through a stochastic process equivalent to a (possibly infinite) linear combination of basis functions. Let $\phi_{l}(x)$ be a basis function for $l=1,\ldots,L$, where $L$ is the number of basis functions. Define $\Phi(x)=(\phi_{1}(x),\ldots,\phi_{L}(x))$ as the $L$-dimensional vector of basis functions, and we model each function as $f_{j}(x)=\beta_{j}^{\top}\Phi(x)$, a linear combination of the pre-specified basis functions $\phi_{l}(x)$. Therefore, for $j=1,\ldots,K$ we want to construct a prior for $\beta_{1},\ldots,\beta_{K}$ such that $K$ functions, $\beta_{1}^{\top}\Phi(x),\ldots,\beta_{K}^{\top}\Phi(x)$ become (nearly) orthogonal. For $j\neq k$, the inner product of $\beta_{j}^{\top}\Phi(x)$ and $\beta_{k}^{\top}\Phi(x)$ is given by
$\int f_{j}(x)f_{k}(x)dx=\int\beta_{j}^{\top}\Phi(x)\beta_{k}^{\top}\Phi(x)dx=\beta_{j}^{\top}\Omega\beta_{k},$
where $\Omega=\int\Phi(x)\Phi(x)^{\top}dx$. Then, the condition that $f_{j}(x)$ and $f_{k}(x)$ are orthogonal is equivalent to
$\beta_{j}^{\top}\Omega\beta_{k}=0\,.$ (1)
We then sequentially define the joint prior of $(\beta_{1},\ldots,\beta_{K})$. First, the prior of $\beta_{1}$ is defined as $\beta_{1}\sim{\mathcal{N}}_{L}(b_{01},B_{01})$ without any restrictions, where $b_{01}$ and $B_{01}$ are hyperparameters and ${\mathcal{N}}_{L}(\mu,\Sigma)$ denotes the $L$-dimensional normal distribution with mean vector $\mu$ and variance-covariance matrix $\Sigma$. Next, the prior of $\beta_{2}$ is defined, conditionally on $\beta_{1}$, as $\beta_{1}^{\top}\Omega\beta_{2}\sim{\mathcal{N}}(0,\tau_{2}^{2})$ and $H_{2}\beta_{2}\sim{\mathcal{N}}_{L}(b_{02},B_{02})$, assuming prior independence between $\beta_{1}^{\top}\Omega\beta_{2}$ and $H_{2}\beta_{2}$, where $b_{02}$, $B_{02}$ and $\tau_{2}^{2}$ are hyperparameters and $H_{2}$ is $(L-1)\times L$ matrix with rank L1L-1 that is independent of $\beta_{1}$. Note that the constraint (1) exactly holds when $\tau_{2}=0$. This conditional prior specification is equivalent to assigning a multivariate normal prior for $A_{2}(\beta_{1})\beta_{2}$ given $\beta_{1}$, where $A_{2}(\beta_{1})=((\beta_{1}^{\top}\Omega)^{\top},H_{2}^{\top})^{\top}$, that is, the first row vector of $A_{2}(\beta_{1})$ is $\beta_{1}^{\top}\Omega$ and the other part is equivalent to $H_{2}$. Hence, the conditional prior of $\beta_{2}$ given $\beta_{1}$ is given by
$\beta_{2}\mid\beta_{1}\sim{\mathcal{N}}_{L}\Big{(}A_{2}(\beta_{1})^{-1}b_{02}^{\ast},A_{2}(\beta_{1})^{-1}{\rm blockdiag}(\tau_{2}^{2},B_{02})\{A_{2}(\beta_{1})^{\top}\}^{-1}\Big{)},$
where $b_{02}^{\ast}=(0,b_{02}^{\top})^{\top}$. Similarly, given $\beta_{1:j}\equiv(\beta_{1},\ldots,\beta_{j})$, the conditional prior for $\beta_{j+1}$ is defined as $\beta_{k}^{\top}\Omega\beta_{j+1}\sim{\mathcal{N}}(0,\tau_{j+1}^{2})$ independently for $k=1,\ldots,j$ and $H_{j+1}\beta_{j+1}\sim{\mathcal{N}}_{L-j}(b_{0,j+1},B_{0,j+1})$, assuming prior independence among $\beta_{k}^{\top}\Omega\beta_{j+1}\ (k=1,\ldots,j)$ and $H_{j+1}\beta_{j+1}$, with hyperparameters $b_{0,j+1}$, $B_{0,j+1}$ and $\tau_{j+1}^{2}$. Here $H_{j+1}$ is an $(L-j)\times L$ matrix that is independent of $\beta_{1:j}$. Again, this prior formulation is equivalent to assigning a multivariate normal prior for $A_{j+1}(\beta_{1:j})\beta_{j+1}$, where $A_{j+1}(\beta_{1:j})=((\beta_{1}^{\top}\Omega)^{\top},\ldots,(\beta_{j}^{\top}\Omega)^{\top},H_{j+1}^{\top})^{\top}$, that is, the first $k$th row vector is $\beta_{k}^{\top}\Omega$ for $k=1,\ldots,j$, and the other part is $H_{j+1}$. Hence, the conditional prior of $\beta_{j+1}$ given $\beta_{1:j}$ is given by
$\beta_{j+1}\mid\beta_{1:j}\sim{\mathcal{N}}_{L}\Big{(}A_{j+1}^{-1}b_{0,j+1}^{\ast},A_{j+1}^{-1}{\rm blockdiag}(\tau_{j+1}^{2}I_{j},B_{0,j+1})(A_{j+1}^{\top})^{-1}\Big{)},$ (2)
for $j=1,\ldots,K-1$, where $A_{j+1}\equiv A_{j+1}(\beta_{1:j})$ and $b_{0,j+1}^{\ast}=(0_{j}^{\top},b_{0,j+1}^{\top})^{\top}$.
The specification of $b_{0,j}$ and $B_{0,j}$ can be arbitrary. For example, if we set $b_{0,j}=0_{L-j}$ and $B_{0,j}=\gamma I_{L-j}$ for some large value of $\gamma$, leading to a diffuse prior. On the other hand, $\tau_{j}^{2}$ controls the exactness of the orthogonal constraints; the constraints are exact when $\tau_{j}^{2}=0$. In many applications, it is important to give nearly exact constraints on $f_{1}(x),\ldots,f_{J}(x)$ for small $J<K$, playing as main principal functions. Hence, it would be better to set smaller values of $\tau_{j}^{2}$ for small value of $j$.
Adaptation of orthogonality constraint
While the proposed prior can be applied with fixed values of $\tau_{k}^{2}$, the specific choice of these hyperparameters can significantly influence the resulting posterior inference. In practice, strict orthogonality (i.e., $\tau_{k}^{2}=0$) may be desirable for leading functions to ensure interpretability, whereas milder constraints (i.e., larger $\tau_{k}^{2}$) for higher-order functions allow more flexibility and help capture finer structures in the data.
To accommodate data-driven adaptation of these constraints, we introduce a hierarchical approach by assigning prior distributions to $\tau_{k}^{2}$. Specifically, we consider two strategies: a global prior and a local prior. In the global prior approach, we assume a common strength of constraint across all components, that is, $\tau_{k}^{2}=\tau^{2}$ for all $k$, and assign an inverse-gamma prior $\tau^{2}\sim\mathrm{IG}(a_{0},b_{0})$ with fixed hyperparameters $(a_{0},b_{0})$. This specification enforces a uniform level of orthogonality and is suitable when the same degree of orthogonality is desired across all functions. In contrast, the local prior approach allows component-specific adaptation by assuming $\tau_{k}^{2}\sim\mathrm{IG}(a_{0},b_{0})$ independently for each $k$. This formulation enables greater flexibility, allowing stronger constraints on leading functions while relaxing constraints on higher-order functions, which can be advantageous when capturing complex variations in the data. Because of such adaptivity of strength of orthogonal constraint, we call the prior adaptively orthogonal prior (AOP).
From a computational perspective, introducing priors for $\tau_{k}^{2}$ only adds an additional inverse-gamma update step in the Gibbs sampler, preserving the overall simplicity of the posterior computation. Regarding the hyperparameters $(a_{0},b_{0})$, we use the default setting $a_{0}=3$ and $b_{0}=2/K^{2}$, which implies that the prior mean of $\tau^{2}$ is $E[\tau^{2}]=1/K^{2}$. This choice provides a weakly informative prior that favors moderate orthogonality (as $\tau^{2}$ is shrunk around a small value), while still allowing sufficient flexibility to adapt to the data.
Properties of the AOP prior
In this section, we investigate properties of the proposed AOP prior and clarify the effect of the hyperparameters. For simplicity, we assume that $\Omega=I_{L}$, that is, the basis functions are orthonormal. Furthermore, we consider the choice of hyperparameter as $b_{0,j}=0_{L-j}$ and $B_{0,j}=\gamma I_{L-j}$ for some $\gamma>0$, and assume that the $(L-j)\times L$ matrix $H_{j+1}$ is orthogonal, that is, $H_{j+1}H_{j+1}^{\top}=I_{L-j}$. Under the setting, we consider the conditional expectation $E[\|\beta_{j+1}\|^{2}|\beta_{1:j}]$, the conditional expectation of the squared norm of the coefficient vector of $f_{j+1}(x)$. It follows that
$A_{j+1}A_{j+1}^{\top}=\left(\begin{array}[]{cc}B_{1:j}^{\top}B_{1:j}&(H_{L-j}B_{1:j})^{\top}\\ H_{L-j}B_{1:j}&I_{L-j}\end{array}\right)$
and
$(A_{j+1}A_{j+1}^{\top})^{-1}=\left(\begin{array}[]{cc}(B_{1:j}^{\top}P_{j}B_{1:j})^{-1}&-(B_{1:j}^{\top}P_{j}B_{1:j})^{-1}B_{1:j}^{\top}H_{L-j}^{\top}\\ -H_{L-j}B_{1:j}(B_{1:j}^{\top}P_{j}B_{1:j})^{-1}&I_{L-j}+H_{L-j}B_{1:j}(B_{1:j}^{\top}P_{j}B_{1:j})^{-1}B_{1:j}^{\top}H_{L-j}^{\top}\end{array}\right),$
where $B_{1:j}=(\beta_{1},\ldots,\beta_{j})$ and $P_{j}=I_{L}-H_{L-j}^{\top}H_{L-j}$. Then, we have
${\rm tr}$ $\{{\rm Var}(\beta_{j+1}|\beta_{1:j})\}$
$={\rm tr}\left\{{\rm blockdiag}(\tau_{j+1}^{2}I_{j},\gamma I_{L-j})(A_{j+1}A_{j+1}^{\top})^{-1}\right\}$
$=\tau_{j+1}^{2}{\rm tr}\{(B_{1:j}^{\top}P_{j}B_{1:j})^{-1}\}+\gamma(L-j)+\gamma\ {\rm tr}\left\{(B_{1:j}^{\top}P_{j}B_{1:j})^{-1}B_{1:j}^{\top}H_{L-j}^{\top}(I_{L}-P_{j})H_{L-j}\right\}.$
Since $E[\beta_{j+1}|\beta_{1:J}]=0_{L}$ from (2), the above expression is equivalent to $E[\|\beta_{j+1}\|^{2}|\beta_{1:j}]$. The conditional expectation of the squared norm (conditional variance) is an increasing function of $\tau_{j}^{2}$, indicating that weaker orthogonal conditions lead to larger conditional variance. Also, under fixed $\tau_{j}^{2}$ and $\gamma$, the conditional variance decreases as $j$ increases, showing that the coefficient $\beta_{j}$ for the $j$th function will be more concentrated around the origin for larger $j$. Such property would be preferable since the $j$th function with large $j$ would not have much contribution.
To visualize the proposed prior, we set $L=K=2$ and $\beta_{1}=(0.5,1)$, and consider the conditonal prior for $\beta_{2}$. In Figure 1, we show the contours of the conditional densities under three choices of the hyperparameters. We observe that when the orthogonality constraint is strong (e.g., $\tau_{2}^{2}=0.01$), the conditional prior distribution is tightly concentrated along the line orthogonal to $\beta_{1}^{\top}\Omega\beta_{2}=0$, effectively enforcing near-orthogonality between the two functions. In contrast, a weaker constraint (e.g., $\tau_{2}^{2}=1$) leads to a broader prior distribution, allowing for more flexibility but reducing orthogonality. Additionally, increasing the variance $B_{02}$ results in a more diffuse prior marginally, while still respecting the constraint level imposed by $\tau_{2}^{2}$. These visualizations highlight how the proposed prior adaptively balances orthogonality and variability through its hyperparameters.
Figure 1: Conditional prior of $\beta_{2}$ given $\beta_{1}$ with $(\tau_{2}^{2},B_{02})=(0.01,1)$ (black), $(1,1)$ (red) and $(0.01,2)$ (blue). The solid line corresponds to the parameter space where the orthogonal condition holds exactly.
Posterior computation
We provide a posterior computation algorithm for the proposed prior. We first note that the joint prior density of $\beta_{1},\ldots,\beta_{K}$ can be expressed as
$\prod_{j=2}^{K}\exp\left\{-\frac{1}{2\tau_{j}^{2}}\sum_{k=1}^{j-1}(\beta_{k}^{\top}\Omega\beta_{j})^{2}\right\}\times\prod_{j=1}^{K}\exp\left\{-\frac{1}{2}(H_{j}\beta_{j}-b_{0j})^{\top}B_{0j}^{-1}(H_{j}\beta_{j}-b_{0j})\right\},$
where $H_{1}=I_{L}$. Then, the full conditional prior of $\beta_{k}$ is proportional to
$\exp\left\{-\frac{1}{2\tau_{k}^{2}}\sum_{j=1}^{k-1}(\beta_{j}^{\top}\Omega\beta_{k})^{2}-\frac{1}{2}\sum_{j=k+1}^{K}\frac{(\beta_{j}^{\top}\Omega\beta_{k})^{2}}{\tau_{j}^{2}}-\frac{1}{2}(H_{k}\beta_{k}-b_{0k})^{\top}B_{0k}^{-1}(H_{k}\beta_{k}-b_{0k})\right\},$
which is a multivariate normal distribution ${\mathcal{N}}_{L}(A_{\beta_{k}}^{-1}B_{\beta_{k}},A_{\beta_{k}}^{-1})$, where
$A_{\beta_{k}}=H_{k}^{\top}B_{0k}^{-1}H_{k}+\frac{1}{\tau_{k}^{2}}\sum_{j=1}^{k-1}\Omega\beta_{j}\beta_{j}^{\top}\Omega+\sum_{j=k+1}^{K}\frac{\Omega\beta_{j}\beta_{j}^{\top}\Omega}{\tau_{j}^{2}},$
$B_{\beta_{k}}=H_{k}^{\top}B_{0k}^{-1}b_{0k}.$
Hence, when the likelihood function for $\beta_{k}$ is Gaussian, we can use a simple Gibbs sampling to generate random draws from the posterior. The specific example will be given in Section 3.
Bayesian functional principal component analysis
We demonstrate the usefulness of the proposed prior in the Bayesian functional principal component analysis (FPCA). Given the functional observations, $X_{1}(t),\ldots,X_{n}(t)$ on $t\in\mathcal{T}$, the Bayesian FPCA assumes the following low-rank model:
$X_{i}(t)=\sum_{k=1}^{K}Z_{ik}f_{k}(t)+{\varepsilon}_{i}(t),\ \ \ \ i=1,\ldots,n,$ (3)
where $f_{k}(t)$ is a principal function common to all the observations, $Z_{ik}$ is an observation-specific random weight for that function and ${\varepsilon}_{i}(t)$ corresponds to an error term. We assume that $Z_{ik}\sim{\mathcal{N}}(0,\lambda_{k})$, where $\lambda_{k}$ is an unknown parameter satisfying $\lambda_{1}>\cdots>\lambda_{K}$. Let $t_{i1},\ldots,t_{im_{i}}$ be a set of observed points for the $i$th observation. By introducing basis function representation for the principal function $f_{k}(x)$, Bayesian inference on FPCA is considered in Suarez and Ghosal (2017) and Nolan et al. (2025), and its robust version is also discussed in Zhang et al. (2025). However, the prior distributions for $f_{k}(x)$ does not guarantee orthogonality each other, which may lead to inefficient low-rank representations. We here introduce the proposed orthogonal prior for a sequence of functions, $f_{1}(x),\ldots,f_{K}(x)$, to obtain (nearly) orthogonal posterior inference on the principal functions.
We perform posterior inference using a Gibbs sampler, which iteratively updates each parameter from its full conditional distribution. The detailed steps are described as follows:
    (Sampling of $\beta_{k}$)   For $k=1,\ldots,K$, draw $\beta_{k}$ from its full conditional distribution, ${\mathcal{N}}_{L}(V_{k}^{-1}U_{k},V_{k}^{-1})$, where
    $\left\{\!\!\begin{array}[]{ll}&\displaystyle V_{k}=H_{k}^{\top}B_{0k}^{-1}H_{k}+\frac{1}{\tau_{k}^{2}}\sum_{j=1}^{k-1}\Omega\beta_{j}\beta_{j}^{\top}\Omega+\sum_{j=k+1}^{K}\frac{\Omega\beta_{j}\beta_{j}^{\top}\Omega}{\tau_{j}^{2}}+\frac{1}{\sigma^{2}}\sum_{i=1}^{n}Z_{ik}^{2}\Phi_{i}\Phi_{i}^{\top},\\ &\displaystyle U_{k}=H_{k}^{\top}B_{0k}^{-1}b_{0k}+\frac{1}{\sigma^{2}}\sum_{i=1}^{n}Z_{ik}\left(X_{i}-\sum_{l\neq k}Z_{il}F_{il}\right).\end{array}\right$.
    Here, $\Phi_{i}=(\phi(t_{i1}),\ldots,\phi(t_{im_{i}}))^{\top}$ is the $m_{i}$-dimensional vector of basis functions evaluated at the observed points, and $F_{il}=(f_{l}(t_{i1}),\ldots,f_{l}(t_{im_{i}}))^{\top}$ denotes the principal function evaluated at the observed points.
    (Sampling of $Z_{ik}$)   For $i=1,\ldots,n$ and $k=1,\ldots,K$, draw $Z_{ik}$ from its full conditional distribution, ${\mathcal{N}}(\mu_{ik}^{\ast},v_{ik}^{\ast})$, where
    $v_{ik}^{\ast}=\left(\frac{F_{ik}^{\top}F_{ik}}{\sigma^{2}}+\frac{1}{\lambda_{k}}\right)^{-1},\quad\mu_{ik}^{\ast}=\frac{v_{ik}^{\ast}F_{ik}^{\top}}{\sigma^{2}}\left(X_{i}-\sum_{l\neq k}Z_{il}F_{il}\right).$
    (Sampling of $\lambda_{k}$)   For $k=1,\ldots,K$, draw $\lambda_{k}$ from its full conditional distribution, ${\rm IG}\left(a_{\lambda}+n/2,\,b_{\lambda}+\sum_{i=1}^{n}Z_{ik}^{2}/2\right)$.
    (Sampling of $\tau_{k}^{2}$)   For $k=1,\ldots,K$, draw $\tau_{k}^{2}$ from its full conditional distribution, $\sim{\rm IG}\left(a_{0}+(k-1)/2,\,b_{0}+\sum_{j=1}^{k-1}(\beta_{j}^{\top}\Omega\beta_{k})^{2}/2\right)$.
    (Sampling of $\sigma^{2}$)   Draw $\sigma^{2}$ from its full conditional distribution, ${\rm IG}(a_{\sigma}+\sum_{i=1}^{n}m_{i}/2,b_{\sigma}+\sum_{i=1}^{n}\sum_{j=1}^{m_{i}}(X_{ij}-\sum_{k=1}^{K}Z_{ik}f_{k}(t_{ij}))^{2}/2)$.
As shown above, all these full conditional distributions are standard distributions, which allows for efficient posterior computation without the need for complex sampling algorithm like rejection steps.
Simulation
We conducted simulation experiments to evaluate the performance of the proposed orthogonality-inducing priors in the context of functional principal component analysis (FPCA) described in Section 3. We considered two scenarios for the underlying true principal functions: (i) Legendre polynomials and (ii) Haar wavelets. For both scenarios, functional observations were generated at $T=30$ equally spaced time points on $[0,1]$, and the number of samples $n$ was set to 50, 100, or 200. In scenario (i), the first three true principal functions were defined as scaled Legendre polynomials, and in scenario (ii), as scaled Haar wavelet functions, described as
$f_{1}(t)$ $=\sqrt{3}\,(2t-1),\ \ \ \ f_{2}(t)=\sqrt{5}\,(6t^{2}-6t+1),\ \ \ \ f_{3}(t)=\sqrt{7}\,(20t^{3}-30t^{2}+12t-1)$
for the Legendre polynomial scenario, and
$f_{1}(t)$ $={\rm I}\Big{(}0\leq t<\frac{1}{2}\Big{)}-{\rm I}\Big{(}\frac{1}{2}\leq t\leq 1\Big{)},\ \ \ \ f_{2}(t)=\sqrt{2}\left\{{\rm I}\Big{(}0\leq t<\frac{1}{4}\Big{)}-{\rm I}\Big{(}\frac{1}{4}\leq t<\frac{1}{2}\Big{)}\right\},$
$\hskip 76.82234ptf_{3}(t)=\sqrt{2}\left\{{\rm I}\Big{(}\frac{1}{2}\leq t<\frac{3}{4}\Big{)}-{\rm I}\Big{(}\frac{3}{4}\leq t\leq 1\Big{)}\right\}.$
for the Haar wavelet scenario, where ${\rm I}(\cdot)$ is an indicator function. The latent scores for the first three components were generated independently from normal distributions with decreasing variances to reflect dominant leading components: $\xi_{1}\sim{\mathcal{N}}(0,1)$, $\xi_{2}\sim{\mathcal{N}}(0,0.7^{2})$, and $\xi_{3}\sim{\mathcal{N}}(0,0.5^{2})$. The observed functional data $Y_{i}(t)$ were then obtained by adding Gaussian noise with standard deviation $\sigma=1$ to the true signal.
For the generated dataset, we applied the proposed adaptive orthogonal prior with a global inverse-gamma prior on $\tau^{2}$ (AOP-G), and local inverse-gamma priors on each $\tau_{k}^{2}$ (AOP-L), where cubic B-splines with an intercept are used for basis functions. For comparison, we adopted a non-orthogonal prior (NO) which uses a standard normal prior for each coefficients, and a non-orthogonal shrinkage prior (NO-S) applying the horseshoe prior (Carvalho et al., 2009) for coefficients of basis functions to reduce the number of irrelevant principal functions. Moreover, we employed nearly mutually orthogonal (NeMO) processes (Matuk et al., 2022). The (largest) number of possible components was set to $K=10$ in all the methods, and the number of basis functions in NO, NO-S, AOP-G and AOP-L was set to $L=12$, noting that the results were not sensitive to the choice of the different value of $L$. For NeMO, the threshold value for orthogonal constraint is set to $10^{-4}$, following the default setting of the R package provided by the author’s GitHub repository. Posterior inference was conducted via MCMC, running for 3,000 iterations after discarding 2,000 burn-in iterations.
We first show the result with a single simulation data under scenario (i). In Figure 2, we visualize the estimated principal functions and heatmaps of their inner products, obtained by “NO” (non-orthogonal) and “AOP-G” (orthogonal) methods. Overall, the results demonstrated that our proposed orthogonal priors successfully enforced near-orthogonality, effectively selected the number of components, and improved reconstruction accuracy while maintaining reasonable uncertainty quantification. In particular, the orthogonal approaches accurately estimated the true number of principal components (three), with the remaining components being automatically shrunk toward zero. This shrinkage behavior is evident in the estimated principal functions, where only the leading three functions show substantial variation, while the others are effectively eliminated. Furthermore, the heatmaps of inner products provide a clear contrast between the methods: under the proposed orthogonal priors, the off-diagonal elements are nearly zero, indicating that the estimated functions are nearly orthogonal. In contrast, the non-orthogonal methods exhibit large off-diagonal inner products, implying strong correlations among components and reducing interpretability. These findings highlight the advantage of orthogonality-inducing priors in achieving interpretable, parsimonious representations in functional principal component analysis.
Figure 2: Estimated principal functions (upper) and heatmaps of inner products of principal functions (lower).
We next evaluate the performance through Monte Carlo simulation studies. For each replication, we evaluated the following two metrics:
    Number of effective components (NC): The number of components with sufficiently large variations, defined as ${\rm NC}=\sum_{k=1}^{K}{\rm I}\big{\{}\int\widehat{f}_{k}^{2}(t)dt>\epsilon\big{\}}$, where $\widehat{f}_{k}(t)\ (k=1,\ldots,K)$ is the posterior mean of the principal functions. We set $\epsilon=0.1$ in this study.
    Orthogonality measure (OG): The sum of absolute off-diagonal inner products among estimated principal functions, defined as ${\rm OG}=\sum_{k=2}^{K}\sum_{j=1}^{k-1}\Big{|}\int\widehat{f}_{j}(t)\widehat{f}_{k}(t)dt\Big{|}$. Note that the smaller values indicate stronger orthogonality.
Tables 1 and 2 summarize the average number of estimated principal components and the orthogonality measures, respectively, based on 200 Monte Carlo replications. Table 1 shows that the non-orthogonal model without shrinkage (NO) consistently overestimates the number of components, returning the maximum value of $K=10$ in all cases. Although the non-orthogonal model with shrinkage (NO-S) reduces the number of estimated components, it still tends to overfit, especially when the sample size is small. In contrast, the adaptive orthogonal prior approaches (AOP-G and AOP-L) provide NC values close to the true number of components, demonstrating the effectiveness of orthogonality-inducing priors in promoting sparsity. On the other hand, NeMO tends to produce NC values larger than the true number of components. Table 2 presents the orthogonality measures based on the sum of absolute off-diagonal inner products. The proposed orthogonal priors yield near-zero values, indicating that the estimated principal functions are nearly orthogonal. On the other hand, both NO and NO-S exhibit large off-diagonal values, especially in small-sample settings, reflecting strong dependencies among components. Also, it is observed that the OG values of NeMo are slightly larger than those of AOP-G and AOP-L. These results confirm that the proposed priors not only improve parsimony but also enhance interpretability by ensuring near-orthogonal decomposition.
Table 1: Average values of estimated number of principal component functions (NC), based on 200 Monte Carlo replications. The standard deviations are given in the parenthesis. True number of components is 3.
Scenario $n$ NO NO-S NeMO AOP-G AOP-L
50 10.0 (0.00) 5.84 (1.63) 7.12 (1.14) 3.31 (0.56) 3.43 (0.59)
1 100 10.0 (0.00) 5.75 (1.53) 7.44 (1.05) 3.14 (0.37) 3.15 (0.39)
200 10.0 (0.07) 6.25 (1.67) 7.05 (1.08) 3.10 (0.36) 3.28 (0.51)
50 10.0 (0.00) 3.30 (0.62) 7.09 (1.17) 3.31 (0.52) 3.32 (0.53)
2 100 10.0 (0.00) 3.09 (0.32) 7.21 (1.25) 3.04 (0.18) 3.02 (0.14)
200 10.0 (0.00) 3.26 (0.48) 7.00 (1.03) 3.00 (0.00) 3.08 (0.26)
Table 2: Average values of the orthogonality measure (OG) of principal functions, based on 200 Monte Carlo replications. The standard deviations are given in the parenthesis.
Scenario $n$ NO NO-S NeMO AOP-G AOP-L
50 16.5 (3.59) 2.45 (1.06) 0.85 (0.21) 0.07 (0.14) 0.07 (0.11)
1 100 22.2 (5.05) 3.11 (1.46) 0.87 (0.19) 0.01 (0.03) 0.04 (0.24)
200 30.9 (8.15) 4.61 (3.11) 0.84 (0.19) 0.01 (0.01) 0.34 (0.67)
50 9.62 (2.22) 2.24 (1.33) 0.63 (0.22) 0.08 (0.16) 0.09 (0.14)
2 100 14.5 (3.66) 2.50 (1.56) 0.65 (0.21) 0.01 (0.04) 0.02 (0.07)
200 20.7 (5.95) 2.92 (2.04) 0.61 (0.16) 0.00 (0.00) 0.12 (0.39)
Data Analysis
We demonstrate the usefulness of the proposed orthogonal prior in Bayesian FPCA, using the population data collected by NTT Docomo Inc., which is the predominant mobile company in Japan. Following the previous studies (Wakayama et al., 2025; Wakayama and Sugasawa, 2024), we focus on the seven special wards of Tokyo metropolitan area and 500-meter mesh grids covering the seven wards. While the original data were provided as three-dimensional arrays, representing locations, time points, and dates, we use the aggregated data consisting of average hourly population counts in February 2019 on weekdays, which results in $n=452$ functional observations with $m_{i}\equiv m=24$ discrete observed points. For $i=1,\ldots,n$, let $Y_{i}^{\ast}(t)$ be the aggregated functional observation, and we use the scaled version $Y_{i}(t)$ defined as $Y_{i}(t)=Y_{i}^{\ast}(t)/\sqrt{m^{-1}\sum_{j=1}^{m}Y_{i}^{\ast}(t_{j})^{2}}$. Figure 3 shows the scaled functional data, which clearly exhibits patterns indicate strong temporal variation, which motivates the use of a low-rank functional representation via FPCA.
Figure 3: Scaled functional observations ($n=452$) of populations in Tokyo area.
We fitted the Bayesian FPCA with four priors, NO, NO-S, NeMO and AOP-G, to these functional data. We adopted the same settings for the prior distributions and tuning parameters as used in Section 4, such as $K=10$ (the largest number of principal functions) and $L=12$ (the number of basis functions). We generated 5,000 iterations and 5,000 burn-in. In Figure 4, we present the posterior distribution of $\tau$ (scale parameter controlling the degree of orthogonality constraint) in AOP-G. These results show that $\tau$ can be successfully learned from the data and its posterior mean is not extremely small, which is a different feature from that of NeMo imposing nearly orthogonal constraint.
Figure 5 displays the posterior means of the principal functions for each method, where only principal functions having variations, $\int\widehat{f}_{k}(t)^{2}dt$, larger than 0.2, are expressed by a solid line. The number of such principal functions are 10 (NO), 6 (NO-S), 5 (NeMO) and 4 (AOP-G). Also, we present inner produces of the estimated principal functions in Figure 6. From these figures, it is observed that the proposed AOP-G method yields interpretable leading components, effectively capturing dominant temporal variations such as morning and evening population peaks. Notably, the orthogonal priors automatically eliminate unnecessary components by shrinking their amplitudes toward zero, resulting in a sparse low-rank representation that clearly highlights the dominant modes of variation. In contrast, the non-orthogonal approaches (NO and NO-S) suffer from overlapping principal components, even when shrinkage prior is applied. While the NeMO method provides more orthogonal principal functions than NO and NO-S, the first principal function is dominant and the other significant principal functions are relatively similar, compared with those of AOP-G. In fact, the orthogonality measures used in Section 4 are 1.72 (NeMO) and 0.18 (AOP-G), which would indicate that AOP-G provides more orthogonal principal functions.
Figure 4: Trace plot of the posterior samples (left) and posterior and prior densities (right) of the scale parameter $\tau$ to control the strength of orthogonality in AOP-G.
To further investigate the usefulness of the proposed methods, we projected the functional observations into two dimensional principal component score, $(Z_{i1},Z_{i2})$, which are shown in Figure 7. The colors in the scatter plots represent two groups obtained via $k$-means clustering on the original functional data (black and red). We observe that the orthogonal approaches, NeMO and AOP-G result in clearer separation of clusters than the non-orthogonal approaches, NO and NO-S. Further, it can be seen that two clusters are successfully separated only by the first principal component function detected by AOP-G. This would exhibit the advantage of AOG-P to extract meaningful principal functions and provide efficient low-rank representations for functional data.
Figure 5: Estimated principal functions, where principal functions with small variations are expressed by dotted grey lines.
Figure 6: The heatmaps of inner products of principal functions.
Figure 7: Scatter plots of the loadings of the first and second principal functions, corresponding to $(Z_{i1},Z_{i2})$ in the FPCA (3), obtained from the four methods.
Concluding Remarks
In this paper, we have proposed a novel class of prior distributions for sequences of orthogonal functions, motivated by applications in functional data analysis and beyond. Our approach introduces a hierarchical construction that sequentially defines function coefficients via conditionally Gaussian distributions with approximate orthogonality constraints. By controlling these constraints through hyperparameters, our framework offers a flexible balance between strict orthogonality and modeling flexibility. Despite these promising results, there remain several directions for future research. While we focused on basis function representations with a finite number of basis, the framework could be potentially extended to Gaussian process models, given the equivalence between GP priors and basis function representations with random coefficients. Such extensions would provide an adaptive version of the constraint prior developed by Matuk et al. (2022).
Acknowledgement
This work is partially supported by JSPS KAKENHI Grant Numbers 24K21420 and 25H00546.